# Strategic Content Creation in the Age of GenAI: To Share or Not to Share?

### v1

@@ B.2 Proofs of Section 3’s Main Results
Lemma 10 suggests that we can restrict attention to the effective strategy space $[0,x_{i}^{\max}]$ for each creator $i\in[n]$, where $x_{i}^{\max}$ is the unique positive root of $\mu x^{\gamma}=c_{i}(x)$. This strategy space is compact and convex.
Step 3: Strict concavity of utility functions.
To establish the strict concavity of $U_{i}$ with respect to $x_{i}$, we compute the second derivative. Applying the product rule and noting that $\frac{\partial\left\|\mathbf{x}_{-i}\right\|}{\partial x_{i}}=0$:
$\frac{\partial^{2}U_{i}}{\partial x_{i}^{2}}=K\left[(\gamma-2)\left\|\mathbf{x}\right\|^{\gamma-3}(\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\left\|\mathbf{x}\right\|^{\gamma-2}\gamma\right]-c_{i}^{\prime\prime}(x_{i}).$
Factoring out $\left\|\mathbf{x}\right\|^{\gamma-3}$:
$\frac{\partial^{2}U_{i}}{\partial x_{i}^{2}}$ $=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma-2)(\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\gamma\left\|\mathbf{x}\right\|\right]-c_{i}^{\prime\prime}(x_{i})$
$=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma^{2}-2\gamma)x_{i}+(\gamma-2)\left\|\mathbf{x}_{-i}\right\|+\gamma(x_{i}+\left\|\mathbf{x}_{-i}\right\|)\right]-c_{i}^{\prime\prime}(x_{i})$
$=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma^{2}-\gamma)x_{i}+(2\gamma-2)\left\|\mathbf{x}_{-i}\right\|\right]-c_{i}^{\prime\prime}(x_{i})$
$=K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right]-c_{i}^{\prime\prime}(x_{i}).$
@@
For $\gamma\in[0,1]$, the term $K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right]$ is non-positive since $\gamma-1\leq 0$, while all other factors are non-negative. Additionally, our assumption of strict convexity implies $c_{i}^{\prime\prime}(x_{i})>0$. Therefore, $\frac{\partial^{2}U_{i}}{\partial x_{i}^{2}}<0$, establishing that $U_{i}$ is strictly concave in $x_{i}$.
Step 4: Monotonicity of the game.
@@
$-\sum_{i=1}^{n}(c_{i}^{\prime}(x_{i}^{\prime})-c_{i}^{\prime}(x_{i}))(x_{i}^{\prime}-x_{i})<0.$
For the first term, we need to analyze $\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})$ in detail. First, we recall the definition of $A_{i}$:
$A_{i}(\mathbf{x})=K\|\mathbf{x}\|^{\gamma-2}\left[\gamma x_{i}+\|\mathbf{x}_{-i}\|\right].$
We need to calculate $A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x})$. By the multivariable mean value theorem, there exists a point $\hat{\mathbf{x}}$ on the line segment between $\mathbf{x}$ and $\mathbf{x}^{\prime}$ such that:
@@
$\frac{\partial A_{i}}{\partial x_{i}}(\mathbf{x})=K\|\mathbf{x}\|^{\gamma-3}(\gamma-1)\left[\gamma x_{i}+2\|\mathbf{x}_{-i}\|\right].$
@@
$=K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)\left[\gamma\hat{x}_{i}+2\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{i}-x_{i})$
$+\sum_{j\neq i}K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)\left[(\gamma-1)\hat{x}_{i}+\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{j}-x_{j}).$
@@
$\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})$ $-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})=\sum_{i=1}^{n}K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)\left[\gamma\hat{x}_{i}+2\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{i}-x_{i})^{2}$
$+\sum_{i=1}^{n}\sum_{j\neq i}K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)\left[(\gamma-1)\hat{x}_{i}+\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{j}-x_{j})(x^{\prime}_{i}-x_{i}).$
Factoring out $K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)$, which is a constant with respect to the summation:
$=K\|\hat{\mathbf{x}}\|^{\gamma-3}(\gamma-1)$ $\bigl{[}\sum_{i=1}^{n}\left[\gamma\hat{x}_{i}+2\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{i}-x_{i})^{2}$
$+\sum_{i=1}^{n}\sum_{j\neq i}\left[(\gamma-1)\hat{x}_{i}+\|\hat{\mathbf{x}}_{-i}\|\right](x^{\prime}_{j}-x_{j})(x^{\prime}_{i}-x_{i})\bigr{]}.$
The expression inside the square brackets matches the structure of $\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{v})$ as defined in Lemma 11.
$\mathcal{Q}_{\gamma}(\mathbf{x},\mathbf{v})=\sum_{i=1}^{n}v_{i}^{2}\left(2\gamma x_{i}+4\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}x_{j}\right)+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}v_{i}v_{j}\left(\gamma(x_{i}+x_{j})+2\sum_{\begin{subarray}{c}k=1\\ k\neq i,j\end{subarray}}^{n}x_{k}\right).$
With a term-by-term comparison, it can be verified that:
$\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})=\frac{1}{2}K(\gamma-1)\|\hat{\mathbf{x}}\|^{\gamma-3}\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{(x^{\prime}-x)}).$
By Lemma 11, we know that $\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{v})\geq 0$ for any $\hat{\mathbf{x}}\geq 0$ and any $\mathbf{v}$. Since $K>0$, $\gamma-1\leq 0$ (as $\gamma\in[0,1]$), and $\|\hat{\mathbf{x}}\|^{\gamma-3}>0$ (because $\|\hat{\mathbf{x}}\|>0$ for non-trivial strategy profiles), we have:
$K(\gamma-1)\|\hat{\mathbf{x}}\|^{\gamma-3}\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{(x^{\prime}-x)})\leq 0.$
@@
Step 5: Existence and uniqueness of equilibrium.
@@
  • Strategy spaces $[0,x_{i}^{\max}]$ are compact and convex.
@@
Proof of Proposition 4.
The proof is given in Appendix F, where it covers a more general form of $Q_{\text{AI}}$. ∎
@@
Step 1: Properties of the ESE profile.
Let $\mathbf{x}^{\star}$ denote the unique ESE whose existence and uniqueness are guaranteed by Proposition 3. By Lemma 12, any rational strategy must involve a strictly positive quality choice, which implies $x_{i}^{\star}>0$ for all $i\in[n]$. This allows us to use Proposition 2 in the next step.
Step 2: Optimality of full sharing under the specified condition.
@@
$\rho>\tau_{i}(x_{i})\coloneqq\frac{x_{i}}{x_{i}+\|\mathbf{x}^{\star}_{-i}\|_{1}+\alpha\|\mathbf{x}^{\star}_{-i}\|_{1}}=\frac{x_{i}}{x_{i}+(1+\alpha)\|\mathbf{x}^{\star}_{-i}\|_{1}}.$
@@
$\frac{\partial\tau_{i}}{\partial x_{i}}=\frac{(1+\alpha)\|\mathbf{x}^{\star}_{-i}\|_{1}}{(x_{i}+(1+\alpha)\|\mathbf{x}^{\star}_{-i}\|_{1})^{2}}>0.$
This confirms that $\tau_{i}(x_{i})$ is strictly increasing in $x_{i}>0$. Recall that Lemma 10 ensures that for any rational quality choice, we must have $x_{i}\leq x_{i}^{\max}$. Since $\tau_{i}(x_{i})$ is strictly increasing in $x_{i}$ and any rational quality choice satisfies $x_{i}\leq x_{i}^{\max}$, the maximum threshold value for creator $i$ occurs at $x_{i}=x_{i}^{\max}$, yielding:
$\max_{x_{i}\in(0,x_{i}^{\max}]}\tau_{i}(x_{i})=\tau_{i}(x_{i}^{\max})=\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\|\mathbf{x}^{\star}_{-i}\|_{1}}.$
@@
$\rho>\max_{i\in[n]}\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\|\mathbf{x}^{\star}_{-i}\|_{1}}.$
@@
Step 3: Verification of FSE properties.
Although this is a direct consequence of the previous steps, we show for completeness that no creator has a profitable deviation from the profile $(\mathbf{x}^{\star},\mathbf{1})$. Consider an arbitrary creator $i$ contemplating a deviation from $(x_{i}^{\star},1)$ to any alternative strategy $(x_{i}^{\prime},s_{i}^{\prime})$.
@@
By Lemma 10, if $(x_{i}^{\prime},s_{i}^{\prime})$ is to be a beneficial deviation, we must have $x_{i}^{\prime}\leq x_{i}^{\max}$. For any such rational quality choice $x_{i}^{\prime}>0$, Step 2 has established that $s_{i}=1$ strictly dominates any other sharing level when $\rho$ satisfies the theorem’s condition. Therefore, regardless of the quality choice $x_{i}^{\prime}$, creator $i$ would maximize utility by choosing $s_{i}=1$ rather than any $s_{i}^{\prime}<1$.
@@
This means that any rational deviation must effectively reduce to a quality-only deviation with full sharing, i.e., $(x_{i}^{\prime},1)$. However, from the definition of an ESE, $x_{i}^{\star}$ is the unique best response to $(\mathbf{x}_{-i}^{\star},\mathbf{1}_{-i})$ among all strategies with $s_{i}=1$. Therefore, no deviation to any $(x_{i}^{\prime},1)$ can be profitable.

### v2

@@ B.2 Proofs of Section 3’s Main Results
Lemma˜13 suggests that we can restrict attention to the effective strategy space $0,x_{i}^{\max}]$ for each creator $i\in[n]$, where $x_{i}^{\max}$ is the unique positive root of $\mu x^{\gamma}=c_{i}(x)$. This strategy space is compact and convex. Additionally, Lemma˜15 ensures that all rational strategies involve strictly positive quality choices, allowing us to focus on the interior of this space.
Step 3: Strict concavity of utility functions
We establish that $U_{i}$ is strictly concave with respect to $x_{i}$ for any fixed $\mathbf{x}_{-i}$ by demonstrating that $R_{i}$ is concave in $x_{i}$, and then using the strict concavity of $c_{i}$ and the fact that the sum of a concave and a strictly concave function is strictly concave. Taking the derivative of $A_{i}$ with respect to $x_{i}$:
$\frac{\partial A_{i}}{\partial x_{i}}$ $=K\left[(\gamma-2)\left\|\mathbf{x}\right\|^{\gamma-3}\left(\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|\right)+\left\|\mathbf{x}\right\|^{\gamma-2}\gamma\right]$
$=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma-2)(\gamma x_{i}+\left\|\mathbf{x}_{-i}\right\|)+\gamma\left\|\mathbf{x}\right\|\right]$
$=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma^{2}-2\gamma)x_{i}+(\gamma-2)\left\|\mathbf{x}_{-i}\right\|+\gamma(x_{i}+\left\|\mathbf{x}_{-i}\right\|)\right]$
$=K\left\|\mathbf{x}\right\|^{\gamma-3}\left[(\gamma^{2}-\gamma)x_{i}+(2\gamma-2)\left\|\mathbf{x}_{-i}\right\|\right]$
$=K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right]$
@@
For $\gamma\in[0,1]$, the term $K(\gamma-1)\left\|\mathbf{x}\right\|^{\gamma-3}\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right]$ is non-positive since $\gamma-1\leq 0$, while all other factors are non-negative. Therefore, $\frac{\partial^{2}R_{i}}{\partial x_{i}^{2}}=\frac{\partial A_{i}}{\partial x_{i}}\leq 0$, establishing that $R_{i}$ is concave in $x_{i}$. Additionally, since $c_{i}$ is strictly convex, $c_{i}$ is strictly concave. Thus, $U_{i}=R_{i}-c_{i}$ is strictly concave in $x_{i}$.
Step 4: Monotonicity of the game
@@
$\sum_{i=1}^{n}(c_{i}^{\prime}(x_{i}^{\prime})-c_{i}^{\prime}(x_{i}))(x_{i}^{\prime}-x_{i})<0.$
For the first term, we need to analyze $\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})$ in detail. We start with $A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x})$. By the multivariable mean value theorem, there exists a point $\hat{\mathbf{x}}$ on the line segment between $\mathbf{x}$ and $\mathbf{x}^{\prime}$ such that:
@@
$\frac{\partial A_{i}}{\partial x_{i}}(\mathbf{x})=K\left\|\mathbf{x}\right\|^{\gamma-3}(\gamma-1)\left[\gamma x_{i}+2\left\|\mathbf{x}_{-i}\right\|\right].$
@@
$=K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)\left[\gamma\hat{x}_{i}+2\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{i}-x_{i})$
$+\sum_{j\neq i}K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)\left[(\gamma-1)\hat{x}_{i}+\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{j}-x_{j}).$
@@
$\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})$ $-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})=\sum_{i=1}^{n}K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)\left[\gamma\hat{x}_{i}+2\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{i}-x_{i})^{2}$
$+\sum_{i=1}^{n}\sum_{j\neq i}K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)\left[(\gamma-1)\hat{x}_{i}+\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{j}-x_{j})(x^{\prime}_{i}-x_{i}).$
Factoring out $K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)$, which is a constant with respect to the summation:
$=K\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}(\gamma-1)$ $\bigl[\sum_{i=1}^{n}\left[\gamma\hat{x}_{i}+2\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{i}-x_{i})^{2}$
$+\sum_{i=1}^{n}\sum_{j\neq i}\left[(\gamma-1)\hat{x}_{i}+\left\|\hat{\mathbf{x}}_{-i}\right\|\right](x^{\prime}_{j}-x_{j})(x^{\prime}_{i}-x_{i})\bigr].$
Next, denote $\mathbf{d}=\mathbf{x}^{\prime}-\mathbf{x}$, and let $B_{ij}=\left((\gamma-1)\hat{x}_{i}+\left\|\hat{\mathbf{x}}_{-i}\right\|\right)$ for $i\neq j$. Notice that the double summation can be rewritten as $\sum_{i\neq j}B_{ij}d_{i}d_{j}$. We now simplify this expression further. The sum $B_{ij}+B_{ji}$ can be expressed as:
$B_{ij}+B_{ji}=(\gamma-1)(\hat{x}_{i}+\hat{x}_{j})+\left\|\hat{\mathbf{x}}_{-i}\right\|+\left\|\hat{\mathbf{x}}_{-j}\right\|=\gamma(\hat{x}_{i}+\hat{x}_{j})+2\sum_{k\neq i,j}\hat{x}_{k}.$
Because the double summation counts each of the pairs $(i,j)$ and $(j,i)$, we can rewrite it as:
$\sum_{i=1}^{n}\sum_{j\neq i}B_{ij}d_{i}d_{j}=\frac{1}{2}\sum_{i=1}^{n}\sum_{j\neq i}(B_{ij}+B_{ji})d_{i}d_{j}=\frac{1}{2}\sum_{i=1}^{n}\sum_{j\neq i}\left[\gamma(\hat{x}_{i}+\hat{x}_{j})+2\sum_{k\neq i,j}\hat{x}_{k}\right]d_{i}d_{j}.$
Thus, we can factor $0.5$ out of both the first expression and the double summation, leading to:
$\sum_{i=1}^{n}(A_{i}(\mathbf{x}^{\prime})-A_{i}(\mathbf{x}))(x_{i}^{\prime}-x_{i})$
$\qquad=\frac{1}{2}K(\gamma-1)\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}\Bigg[\sum_{i=1}^{n}d_{i}^{2}\left(2\gamma\hat{x}_{i}+4\left\|\hat{\mathbf{x}}_{-i}\right\|\right)+\sum_{i=1}^{n}\sum_{j\neq i}d_{i}d_{j}\left(\gamma(\hat{x}_{i}+\hat{x}_{j})+2\sum_{k\neq i,j}\hat{x}_{k}\right)\Bigg].$
The expression inside the square brackets matches the structure of $\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{d})$ as defined in Lemma˜14. By Lemma˜14, we know that $\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{d})\geq 0$ for any $\hat{\mathbf{x}}\geq 0$ and any $\mathbf{d}$. Since $K>0$, $\gamma-1\leq 0$ (as $\gamma\in[0,1]$), and $\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}>0$ (because $\left\|\hat{\mathbf{x}}\right\|>0$ for non-trivial strategy profiles), we have:
$K(\gamma-1)\left\|\hat{\mathbf{x}}\right\|^{\gamma-3}\mathcal{Q}_{\gamma}(\hat{\mathbf{x}},\mathbf{(x^{\prime}-x)})\leq 0.$
@@
Step 5: Existence and uniqueness of equilibrium
@@
  • Strategy spaces $0,x_{i}^{\max}]$ are compact and convex.
@@
Proof of Proposition˜4.
We prove the result for the more general GenAI quality function $Q_{\text{AI}}=\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}$ with $\beta\in(0,1]$; the linear case $\beta=1$ used in the main text follows as a special case. The proof proceeds in two parts: we first establish the existence and uniqueness of an FSE, and then show that no other PNE exists.
Existence and uniqueness of FSE
We begin by showing that under $\rho=1$, full sharing is the unique optimal choice for any creator with positive quality. Fix a creator $i$ with $x_{i}>0$, and suppose $\left\|\mathbf{x}_{-i}\right\|>0$. Excluding the cost term $c_{i}(x_{i})$, which is independent of $s_{i}$, creator $i$’s utility as a function of $s_{i}$ can be written as
$\widehat{U}_{i}(s_{i})=T(\mathbf{x})\left[\frac{x_{i}+\alpha x_{i}s_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-1}}{\left\|\mathbf{x}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}}\right].$
Differentiating with respect to $s_{i}$ and simplifying yields
$\frac{d\widehat{U}_{i}}{ds_{i}}=T(\mathbf{x})\cdot\frac{\alpha x_{i}(\mathbf{x}^{\top}\mathbf{s})^{\beta-2}}{(\left\|\mathbf{x}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta})^{2}}\left\{\beta x_{i}s_{i}\left\|\mathbf{x}_{-i}\right\|+\mathbf{x}^{\top}_{-i}\mathbf{s}_{-i}\left[(1-\beta)x_{i}+\left\|\mathbf{x}_{-i}\right\|+\alpha(\mathbf{x}^{\top}\mathbf{s})^{\beta}\right]\right\}.$
Under our assumptions ($\alpha>0$, $x_{i}>0$, $\left\|\mathbf{x}_{-i}\right\|>0$), the fraction is strictly positive for all $s_{i}\in(0,1)$. Moreover, the first term in braces is strictly positive, and the second is non-negative. Hence $\frac{d\widehat{U}_{i}}{ds_{i}}>0$ on $(0,1)$, implying that $\widehat{U}_{i}$ is strictly increasing, and therefore $s_{i}=1$ is the unique optimum.
Next, we show that under $\rho=1$ and full sharing, the utility function simplifies to a form that does not depend on $\beta$. When $\mathbf{s}=\mathbf{1}$, we have $\mathbf{x}^{\top}\mathbf{s}=\left\|\mathbf{x}\right\|$ and $f_{i,1}(\mathbf{x},\mathbf{1})=\frac{x_{i}}{\left\|\mathbf{x}\right\|}$. Substituting into the general utility expression:
$U_{i}(\mathbf{x},\mathbf{1};f_{1})=\mu\left\|\mathbf{x}\right\|^{\gamma}\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta}}+\mu\left\|\mathbf{x}\right\|^{\gamma}\cdot\frac{\alpha\left\|\mathbf{x}\right\|^{\beta}}{\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta}}\cdot\frac{x_{i}}{\left\|\mathbf{x}\right\|}-c_{i}(x_{i}).$
Combining the revenue terms over a common denominator and factoring, we obtain
$U_{i}(\mathbf{x},\mathbf{1};f_{1})=\frac{\mu\left\|\mathbf{x}\right\|^{\gamma}x_{i}(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})}{(\left\|\mathbf{x}\right\|+\alpha\left\|\mathbf{x}\right\|^{\beta})\left\|\mathbf{x}\right\|}-c_{i}(x_{i})=\mu\left\|\mathbf{x}\right\|^{\gamma-1}x_{i}-c_{i}(x_{i}).$
This simplified form is identical to that analyzed in the proof of Lemma˜3 for $\beta=1$. Since that analysis depends only on the structure $\mu\left\|\mathbf{x}\right\|^{\gamma-1}x_{i}-c_{i}(x_{i})$ and not on $\beta$, we conclude that the enforced sharing game under $\rho=1$ admits a unique ESE, which we denote by $\mathbf{x}^{\star}$.
To establish that $(\mathbf{x}^{\star},\mathbf{1})$ is the unique FSE, consider any deviation by creator $i$ to $(x_{i}^{\prime},s_{i}^{\prime})$. By a straightforward extension of Lemma˜15 to general $\beta$, we have $x_{j}^{\star}>0$ for all $j$, ensuring $\left\|\mathbf{x}^{\star}_{-i}\right\|>0$. By the analysis above, any profitable deviation must have $s_{i}^{\prime}=1$. But then the best response in quality is $x_{i}^{\prime}=x_{i}^{\star}$, since $\mathbf{x}^{\star}$ is an ESE. Thus $(\mathbf{x}^{\star},\mathbf{1})$ is an FSE, and uniqueness follows from the uniqueness of the ESE.
No other PNE exists
Suppose $(\mathbf{x}^{\prime},\mathbf{s}^{\prime})$ is an arbitrary PNE. By the extension of Lemma˜15, we must have $x_{i}^{\prime}>0$ for all $i\in[n]$. Since $n\geq 2$, this implies $\left\|\mathbf{x}^{\prime}_{-i}\right\|>0$ for each creator $i$. By the optimality of full sharing established above, we must have $\mathbf{s}^{\prime}=\mathbf{1}$. But then the quality profile $\mathbf{x}^{\prime}$ must constitute an ESE, and by uniqueness, $\mathbf{x}^{\prime}=\mathbf{x}^{\star}$. Therefore, $(\mathbf{x}^{\star},\mathbf{1})$ is the unique PNE. ∎
@@
Step 1: Properties of the ESE profile
Recall that the existence and uniqueness of the ESE profile $\mathcal{X}(\rho)$ are guaranteed by Lemma˜3. By Lemma˜15, any rational strategy must involve a strictly positive quality choice, which implies $\mathcal{X}_{i}(\rho)>0$ for all $i\in[n]$. This allows us to use Lemma˜2 in the next step.
Step 2: Optimality of full sharing under the specified condition
@@
$\rho>\tau_{i}(x_{i})\coloneqq\frac{x_{i}}{x_{i}+\left\|\mathcal{X}_{-i}(\rho)\right\|+\alpha\left\|\mathcal{X}_{-i}(\rho)\right\|}=\frac{x_{i}}{x_{i}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}.$
@@
$\frac{\partial\tau_{i}}{\partial x_{i}}=\frac{(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}{(x_{i}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|)^{2}}>0.$
This confirms that $\tau_{i}(x_{i})$ is strictly increasing in $x_{i}>0$. Recall that Lemma˜13 ensures that for any rational quality choice, we must have $x_{i}\leq x_{i}^{\max}$. Since $\tau_{i}(x_{i})$ is strictly increasing in $x_{i}$ and any rational quality choice satisfies $x_{i}\leq x_{i}^{\max}$, the maximum threshold value for creator $i$ occurs at $x_{i}=x_{i}^{\max}$, yielding:
$\max_{x_{i}\in(0,x_{i}^{\max}]}\tau_{i}(x_{i})=\tau_{i}(x_{i}^{\max})=\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}.$
@@
$\rho>\max_{i\in[n]}\frac{x_{i}^{\max}}{x_{i}^{\max}+(1+\alpha)\left\|\mathcal{X}_{-i}(\rho)\right\|}.$
@@
Step 3: Verification of FSE properties
Consider an arbitrary creator $i$ contemplating a deviation from $(\mathcal{X}_{i}(\rho),1)$ to any alternative strategy $(x_{i}^{\prime},s_{i}^{\prime})$.
@@
By Lemma˜13, if $(x_{i}^{\prime},s_{i}^{\prime})$ is to be a beneficial deviation, we must have $x_{i}^{\prime}\leq x_{i}^{\max}$. For any such rational quality choice $x_{i}^{\prime}>0$, Step 2 has established that $s_{i}=1$ strictly dominates any other sharing level when $\rho$ satisfies the theorem’s condition. Therefore, regardless of the quality choice $x_{i}^{\prime}$, creator $i$ would maximize utility by choosing $s_{i}=1$ rather than any $s_{i}^{\prime}<1$.
@@
This means that any rational deviation must effectively reduce to a quality-only deviation with full sharing, i.e., $(x_{i}^{\prime},1)$. However, from the definition of an ESE, $\mathcal{X}_{i}(\rho)$ is the unique best response to $(\mathbf{x}_{-i}^{\star},\mathbf{1}_{-i})$ among all strategies with $s_{i}=1$. Therefore, no deviation to any $(x_{i}^{\prime},1)$ can be profitable.